# Active particles in tunable crowded environments

### v1

@@ Abstract
Active particles affect their environment as much as the environment affects their active motion. Here, we present an experimental system where both can be simultaneously adjusted in situ using an external AC electric field. The environment consists in a two-dimensional bath of colloidal silica particles, whereas the active particles are gold-coated Janus spheres. As the electric field orthogonal to the planar layer increases, the former become stiffer and the latter become faster. The active motion evolves from a viscous-like to a viscoelastic-like behavior, with the reorientation frequency increasing with the particle speed. This effect culminates in the spontaneous chiralization of particle trajectories. We demonstrate that self-sustained reorientations arise from local compressions and interaction asymmetries, revealing a general particle-level mechanism where changes in the mechanical properties of the environment reshape active trajectories.
@@ 1 Introduction
The motion of living and synthetic microswimmers in viscoelastic fluids (e.g. polymer solutions or cytoskeletal fluids) is another paradigmatic example of feedback between activity and environmental properties. In stark contrast to Newtonian fluids, where the persistence time ($\tau$) and swimming speed ($v$) of active particles are typically independent, memory effects associated with viscoelasticity induce a strong coupling between these quantities [20, 21, 22, 23]. Small velocity changes can lead to a decrease (or increase) of orders of magnitude in the persistence time [22], with important consequences for self-transport, self-organization and active flows. In extreme cases, viscoelastic fluids can even break the polar symmetry of the active particles and stabilize chiral helical trajectories [20] and rotating vortices [24]. Despite these dramatic changes and the abundance of non-Newtonian fluids in nature, viscoelastic effects in active matter are largely unexplored and have not yet been harnessed to adjust in situ the motion of active particles.
@@
Since our goal is to tune, reversibly and simultaneously, both the activity of the Janus particles and the stiffness of the colloidal matrix, a careful selection and reproducible preparation of the experimental system are paramount. We opted for gold-capped Janus silica particles actuated by AC electric fields in a bath of passive silica spheres. As compared to other Janus swimmers (e.g. catalytic), AC-driven particles are particularly versatile: they do not ‘get stuck’ in stiff environments and their propulsion direction and speed can be tuned through the frequency and magnitude of the applied electric field [40, 41]. Passive silica spheres also show a frequency dependent transition from attractive to repulsive interactions, each of which increases with increasing field strength [42]. Here, we chose the AC frequency to be 2020 $\rm kHz$, which assures (1) powerful propulsion of the active particles with the gold-capped side at their front, (2) continuously variable long-range repulsion between silica surfaces (i.e. between passive spheres as well as between passive particles and the back hemispheres of the active colloids) and (3) a short-range repulsive interaction between the passive spheres and the gold cap of the Janus particles. As we increase the strength of the orthogonal AC electric field, the colloidal background freezes and becomes stiffer (Fig. 1(a) $\rightarrow$ Fig. 1(b)). Upon this quenching, we witness pronounced viscoelastic effects so that the persistence time $\rm\tau$ of the active particles drops by more than one order of magnitude (compare for example the red trajectories in Fig. 1(a) and Fig. 1(b), or in the corresponding Supplementary Videos S1 and S2). In particular, $\rm\tau$ transitions from being independent of the mean swimming velocity $v$ to scaling as $\tau\propto v^{-1}$. Moreover, for sufficiently large propulsion speeds and silica-silica interactions, we observe for the first time the emergence of chiral helical swimming in a colloidal bath composed of non-chiral objects, in analogy to what has been reported in viscoelastic polymer solutions, but with a different underlying reason [20].
@@
The particle-resolved approach and flexible experimental system allow to identify clear ingredients leading to helical motion and reshape, reversibly and in situ, the active motion from viscous-like ($\tau=\rm const$) to viscoelastic-like ($\tau=\tau(v)$). The mechanism behind the faster reorientation of active particles in stiffer environments is general and based on an interaction asymmetry of the active particle with the discrete colloidal matrix, a feature which our Janus particles share with a large number of other microswimmers, in particular also metal capped catalytic swimmers. The dipolar repulsion acts in a non-central way, i.e. the stronger interaction with the silica side naturally leads to the creation of a lever arm that generates a torque. To bring this mechanism to work, long-ranged interactions and a compressible matrix are needed. The spontaneous chiralization then occurs as a sustained symmetry breaking initially caused by an orientational fluctuation. These findings demonstrate the ability to regulate the reorientation dynamics of active particles through mechanical changes in its surroundings.
@@
The manuscript is structured as follows. In Section 2, we first discuss the preparation of colloidal monolayers where the inter-particle interactions can be tuned using external AC electric fields. We then describe the two-dimensional motion of active particles throughout colloidal monolayers in term of their swimming velocity and persistence time. Here, the electric field determines not only the interactions between the passive particles of the monolayer, but also the propulsion speed of the active beads. In Section 3, we present a final discussion on our results. Section 4 contains the experimental and numerical protocols.
@@ 2.1 Tunable colloidal environments
Figure 2: Tuning the structure of the environment. (a) Experimental (red solid lines) and numerical (black dotted lines) pair correlation functions $\rm g(r)$ for colloidal monolayers with packing fraction $\phi=0.48$ subjected to electric fields of different magnitude (as shown in the graph). The experimental data are an average over $1000$ independent realizations. (b) Mean absolute value of the hexagonal order parameter $|\Psi_{6}|$ as a function of the applied electric field. Experiments and simulations are denoted by solid and empty symbols, respectively. The dotted line connecting the numerical results is a guide for the eye. (c-e) Maps of the phase $\varphi$ of $\rm\Psi_{6,n}$ for experiments at (c) $\rm E=42$ $\rm V/mm$, (d) $\rm E=75$ $\rm V/mm$ and (e) $\rm E=108$ $\rm V/mm$. The color indicates the orientation of the crystalline domain, as shown below in the panel. The scale bar is 2020 $\rm\mu m$.
@@
In simulations, we model the background matrix as a two-dimensional system of interacting Brownian particles, where hydrodynamics interactions are negligible and the solvent simply induces a thermal noise [44]. The particles interact via volume exclusion, modeled as a WCA potential [45], and a long-range dipole-dipole repulsion mimicking the dipolar interactions between $\rm SiO_{2}$ microparticles at $\rm E>0$. The strength of these potentials, $\epsilon$ and $\rm K$ respectively, is determined by matching the peaks of the pair correlation functions $\rm g(r)$ with experiments for every value of $\rm E$ and $\rm\phi$ (see Fig. 2(a), Fig. S1 and Table S1). Further experimental and numerical details are reported in Sec. 4.
Both in simulations and experiments, the long-range repulsive interactions induced by the external electric field promote the shift of the freezing line in the two-dimensional phase diagram towards lower packing fractions. Features that are characteristic of solid structures (e.g. hexagonal crystalline domains) appear at smaller $\phi$, when the original microstructure (at $\rm E=0$) was in the fluid phase. The red curves in Fig. 2(a) show for example the pair correlation functions, $\rm g(r)$, for a monolayer at $\phi=0.48$ subjected to three different electric fields. As $\rm E$ increases, the peaks of $\rm g(r)$ become more pronounced until, at $\rm E=108$ $\rm V/mm$, the second peak between $\rm 4R_{p}$ and $\rm 6R_{p}$ splits, as reported for hexagonal structures. To quantify the appearance of solid domains, we consider the absolute value of the orientational hexagonal order parameter per particle:
@@
Figure 3: Quenching colloidal environments by changing the electric field amplitude. (a) Mean hexagonal bond order parameter $\rm\left|\Psi_{6}\right|$ plotted as a function of the electric-field strength $\rm E$ for colloidal monolayers with packing fraction $\phi=0.15$ (green), $\phi=0.35$ (blue), $\phi=0.48$ (red) and $\phi=0.64$ (orange). The filled symbols are obtained by experiments while the empty symbols (connected by dotted lines) correspond to numerical results. (b) Experimental dynamical orientational correlation $\rm g_{6}(\Delta t)$ as a function of the normalized delay time $\rm\Delta t/\tau_{B}$ for different packing fractions and electric fields. (a) and (b) share the same legend with colors corresponding to different packing fractions $\phi$. Note that in (b) we only report the curves at $\rm E=48$ $\rm V/mm$, $\rm E=75$ $\rm V/mm$ and $\rm E=108$ $\rm V/mm$ for clarity. At all $\phi$, the decay of $\rm g_{6}(\Delta t)$ is slower for larger electric fields (as schematically indicated by the arrows).
The structural changes induced by the electric field are observed for a broad range of packing fractions, suggesting the generality of the observed quenching mechanism. These results are shown in Fig. 3(a), where $\rm\left|\Psi_{6}\right|$ is reported as a function of the electric field $\rm E$ for four different values of $\phi$. The corresponding pair correlation functions $\rm g(r)$ and phases $\varphi$ of $\rm\Psi_{6,n}$ are shown in Fig. S1 and Fig. S2, respectively. In all instances, the global bond order parameter $\rm\left|\Psi_{6}\right|$ increases with $\rm E$, although we observe only a weak increase for large values of $\rm E$ at $\phi=0.15$ (green data) and a plateau in the densest environment ($\phi=0.64$, orange data). Indeed, in the small packing fraction regime, the smaller is the density, the larger is the increase of the dipole interaction strength needed to induce significant structural changes. By contrast, at high packing fractions, the observed plateau implies that the system has already reached the almost-close packing regime characterized by a defects-poor configuration. As a further confirmation, we illustrate in Fig. 3(b) the dynamical orientational correlation $\rm g_{6}(\Delta t)=\langle\Psi^{*}_{6,n}(\Delta t)\Psi_{6,n}(0)\rangle$ as a function of the delay time $\rm\Delta t/\tau_{B}$ normalized by Brownian diffusion time $\rm\tau_{B}=(6\pi\eta_{0}R_{p}^{3})/(k_{B}T)$, where $\eta_{0}$ is the water viscosity and $\rm k_{B}T$ is the thermal energy (see also Fig. S3 for a comparison with numerical results). At $\phi=0.64$ (orange data), $\rm g_{6}(\Delta t)$ shows no significant decay within our experimental timescale, suggesting the presence of defect-free structures. At smaller packing fractions, $\rm g_{6}(\Delta t)$ decays at a rate that depends on both $\rm E$ and $\phi$. Interestingly, similar decays are recovered with different packing fractions and electric fields, in agreement with the measurements of $\rm\left|\Psi_{6}\right|$ (compare for instance the red and blue data in Fig. 3(a)). This demonstrates that the electric field can be used as a tool to create optimally ordered configurations without changing the packing fraction.
@@ 2.2 Active particles in tunable colloidal environments
We then prepare the same colloidal environments as in Figs. 2-3 and add a very small number of self-propelling particles such that the interactions between them can be neglected. The active particles behave as ‘snowploughs’ making their way through the two-dimensional crowded environment (see Supplementary Videos S1 and S2). Even though they do not sense local defects and grain boundaries due to the large size (compared to the one of the bath passive particles), their active dynamics is strongly coupled to the monolayer’s mechanical properties (see Fig. 1a and Fig. 1b). In simulations, this behavior is reproduced through a 2D active Brownian particle dynamics where the active particle moves in the direction identified by the outward vector normal to the $\rm Au$ hemisphere. As for passive colloids, hydrodynamics interactions are neglected and the solvent induces translational and rotational noise on the particle position and orientation [49]. The active particle interacts with passive colloids via volume exclusion and repulsive dipolar forces. Volume exclusion is obtained from a WCA potential accounting for the larger size of the active colloid, whereas dipolar forces stem from the $\rm SiO_{2}$ back hemisphere of the active particle. As such, although active-passive dipolar interactions follow the $\rm~r^{-3}$ scaling (as in the case of uncoated passive particles), the force is not applied on the particle center. Instead, it acts the center of the SiO2 hemisphere, due to the fact that the gold cap and the passive particles do not exhibit dipolar mutual interactions. Hence, the orientational dynamics is not purely governed by rotational noise but it is additionally susceptible to local torques. In the following, we analyze the active trajectories by extracting the mean swimming velocity $v$ and persistence time $\tau$. The same quantities measured at $\phi=0$ (without the colloidal environment) are denoted as $v_{0}$ and $\tau_{0}$. As in the previous Section, we first describe the main results at $\phi=0.48$ and then show the generality of our findings at other packing fractions. Further information about the particle synthesis, self-propulsion method, numerical simulations and theoretical model can be found in Sec. 4.
Figure 4: Swimming in crowded environments. (a) Mean swimming velocities (black) $v_{0}$ and (red) $v$ plotted as a function of the electric-field strength squared $\rm E^{2}$ for active particles swimming ‘freely’ ($\phi=0$) or cruising in colloidal environments of packing fraction $\phi=0.48$. The solid lines are linear fits crossing the origin. The error bars are calculated as standard deviations. (b) Ratio between the effective viscosity $\rm\eta_{eff}$ of the colloidal environment and the water viscosity $\rm\eta_{0}$ calculated as described in the text. (c-d) Normalized histograms of the instantaneous speed $\left|\dot{\mathbf{x}}\right|$ under an applied electric field (c) $\rm E=42$ $\rm V/mm$ and (d) $\rm E=108$ $\rm V/mm$.
@@
Nonetheless, a description based on an average swimming velocity of the self-propelling particles and an effective viscosity of the colloidal environment does not take into account the microstructure of the monolayer. A closer look at the instantaneous velocities sometimes reveals large fluctuations around their mean value, where $\left|\dot{\mathbf{x}}\right|$ drops to nearly zero before increasing again. This local and transient caging is due to the microstructure of the surrounding monolayer as well as the strength of interaction between its building blocks [26, 50]; it is qualitatively illustrated in the Supplementary Video S6 and pinpointed in Fig. 4(c-d) by plotting the probability distribution ${\rm p}(\left|\dot{\mathbf{x}}\right|)$. For small applied electric fields ($\rm E=42$ $\rm V/mm$, Fig. 4(c)), the dipolar interactions are weak and the active particle can navigate through the environment owing to the high mobility of the uncoated $\rm SiO_{2}$ microspheres. The corresponding distribution $\rm p(\left|\dot{\mathbf{x}}\right|)$ is therefore similar to those of free active particles (see also Figure S5, gray histograms). At higher electric fields ($\rm E=108$ $\rm V/mm$, Fig. 4(d)), $\rm p(\left|\dot{\mathbf{x}}\right|)$ shows a peak at $\left|\dot{\mathbf{x}}\right|\sim 0$ followed by a distribution at higher speeds; the active motion becomes intermittent and particles are occasionally caged by the surrounding microstructure. Velocity distributions for intermediate values of $\rm E$ are reported in Fig. S5 and show a consistent transition towards a bimodal distribution, as $\rm E$ increases. We report no caging at smaller packing fractions ($\phi=0.35$ and $\phi=0.15$). The same qualitative behaviour is also observed in numerical simulations (Fig. S5), although caging occurs therein more often. The discrepancy is likely due to the fact that the numerical system is perfectly two dimensional, whereas the experimental active particles protrude above the colloidal monolayer because of their larger size (see Fig. 8 in Sec. 4).
Figure 5: Reorienting in crowded environments. (a-c) Trajectories of active particles cruising in colloidal environments of packing fraction $\phi=0.48$ under applied electric fields (a) $\rm E=42$ $\rm V/mm$, (b) $\rm E=75$ $\rm V/mm$ and (c) $\rm E=108$ $\rm V/mm$. The scale bar is ‘dynamic’ and corresponds to $\left(v\tau_{\rm R}\right)$ for all trajectories, where $v$ is the average swimming velocity and $\rm\tau_{R}$ is the rotational Brownian time (defined in the text). The color indicates the time: from $\rm t=0$ (white) to $\rm t=500$ $\rm s$ (red). (d) Time autocorrelation functions of $\dot{\mathbf{x}}$ for the trajectories shown in (a-c). The solid lines are linear fits. (e) Mean persistence times $\tau$ (red symbols, $\phi=0.48$) and $\tau_{0}$ (black symbols, $\phi=0$) plotted against $v$. The filled symbols are obtained from experiments while empty symbols (linked by the dotted line) are from numerical simulations matching the experimental values of $v$. The error bars are calculated as standard deviations.
@@
An increase of effective viscosity due to larger $\phi$ or $\rm E$ should also imply a slowdown of rotational Brownian motion of the active particles in the $xy$-plane and, consequently, a higher persistence time of their directed motion (i.e. a larger $\tau$). The data in Figure 5 suggest the opposite. Fig. 5(a-c) shows three typical active trajectories at fixed $\phi=0.48$ and different applied electric fields. As $\rm E$ increases from $42$ $\rm V/mm$ (Fig. 5(a)) to $108$ $\rm V/mm$ (Fig. 5(c)), the particles not only become faster, but also change their swimming direction more frequently, as shown by comparing the typical straight path of each trajectories to the persistence length $(v\tau_{\rm R})$ expected as a result of free rotational diffusion with timescale $\rm\tau_{R}$. We extract the characteristic persistence times $\tau$ and $\tau_{0}$ from the linear fit (Fig. 5(d), black lines) of the time autocorrelation function of the instantaneous velocity vector $\rm\left\langle\dot{\mathbf{x}}(\Delta t)\cdot\dot{\mathbf{x}}(0)\right\rangle$ (Figure 5(d), red circles), as described in Sec. 4. Faster decays indicate smaller persistence times. At $\phi=0.48$, the velocity autocorrelation decays faster for larger applied electric fields. This corresponds to mean persistence times that become more than one order of magnitude smaller as $\rm E$ goes from $\rm E=42$ $\rm V/mm$ to $\rm E=108$ $\rm V/mm$, i.e. as the mean swimming velocity increases (Fig. 5(e), red symbols). This is in stark contrast to the motion of ‘free’ active particles ($\phi=0$) where $\tau_{0}$ remains approximately constant at all applied electric fields and swimming velocities (see black symbols in Fig. 5(e) as well as Fig. S4). The measured persistence times are also in agreement with numerical simulations (Fig. 5(e), empty red symbols) that are set to match the mean swimming velocity $v$ of the active particles and the pair correlation function $\rm g(r)$ of the colloidal bath.
Figure 6: Self-sustaining reorientation. (a-b) Sketch of the reorientation mechanism, as modeled in numerical simulations, causing the drop of the persistence time of the active particle in the colloidal environments. (a) An active particle moving with the $\rm Au$ cap heading produces a swimming force $f_{\rm s}\mathbf{\hat{n}}$ in the same direction. (b) Upon rotation by an angle $\Delta\theta$, the colloidal environment is compressed leading to a torque in the same direction due to the dipolar repulsion between passive particles and the silica back hemisphere of the active colloid. (c-d) Voronoi tesselation of the environments (c) before and (d) after a rotation. The color code indicates the local packing, i.e. the inverse area occupied by the Voronoi cell. Upon rotation, the local density distribution of the matrix becomes asymmetric.

### v2

# Active particles in tunable compressible environments

@@ Abstract
Active particles affect their environment as much as the environment affects their active motion. Here, we present an experimental system where both can be simultaneously adjusted in situ using an external AC electric field. The environment consists in a two-dimensional bath of colloidal silica particles, whereas the active particles are gold-coated Janus spheres. As the electric field orthogonal to the planar layer increases, the former become stiffer and the latter become faster. The active trajectories exhibit enhanced rotational motion where the reorientation frequency increases with the particle speed, an effect that culminates in a chiral active motion. We demonstrate that self-sustained reorientations arise from local compressions and interaction asymmetries, revealing a general particle-level mechanism where changes in the mechanical properties of the environment reshape active trajectories.
@@ 1 Introduction
The motion of microswimmers in complex fluids is another paradigmatic example of feedback between activity and environmental properties, with important consequences for self-transport, self-organization and active flows [20, 21]. In particular, recent experiments and simulations reported a dramatic decrease of persistence time [22, 23, 24] and even a break of polar symmetry [25, 26, 27] in polymer solutions and nanostructured media. Despite the abundance of complex fluids in nature, these feedback mechanisms in active matter remain largely unexplored and have not yet been harnessed to adjust in situ the motion of active particles.
@@
Since our goal is to tune, reversibly and simultaneously, both the activity of the Janus particles and the stiffness of the colloidal matrix, a careful selection and reproducible preparation of the experimental system are paramount. We opted for gold-capped Janus silica particles actuated by AC electric fields in a bath of passive silica spheres. As compared to other Janus swimmers (e.g. catalytic), AC-driven particles are particularly versatile: they do not ‘get stuck’ in stiff environments and their propulsion direction and speed can be tuned through the frequency and magnitude of the applied electric field [43, 44]. Passive silica spheres also show a frequency dependent transition from attractive to repulsive interactions, each of which increases with increasing field strength [45]. Here, we chose the AC frequency to be 2020 $\rm kHz$, which assures (1) powerful propulsion of the active particles with the gold-capped side at their front, (2) continuously variable long-range repulsion between silica surfaces (i.e. between passive spheres as well as between passive particles and the back hemispheres of the active colloids) and (3) a short-range repulsive interaction between the passive spheres and the gold cap of the Janus particles. As we increase the strength of the orthogonal AC electric field, the colloidal background freezes due to stronger repulsive interactions whereas the Janus particles become faster. Upon this quenching, the persistence time $\rm\tau$ of the active particles drops by more than one order of magnitude (compare for example the red trajectories in Fig. 1(a) and Fig. 1(b), or in the corresponding Supplementary Videos S1 and S2). In particular, $\rm\tau$ transitions from being independent of the mean swimming velocity $v$ to scaling as $\tau\propto v^{-1}$. Moreover, for sufficiently large propulsion speeds and silica-silica interactions, we observe for the first time the emergence of chiral helical swimming in a bath composed of non-chiral microspheres.
@@
The particle-resolved approach and flexible experimental system allow to identify clear ingredients leading to helical motion and reshape, reversibly and in situ, the rotational characteristics of the active motion. The mechanism behind the faster reorientation is general and based on an interaction asymmetry of the active particle with the discrete colloidal matrix, a feature which our Janus particles share with a large number of other microswimmers. The dipolar repulsion acts in a non-central way, i.e. the stronger interaction with the silica side naturally leads to the creation of a lever arm that generates a torque. To bring this mechanism to work, long-ranged interactions and a compressible matrix are needed. The spontaneous chiralization then occurs as a sustained symmetry breaking initially caused by an orientational fluctuation. These findings demonstrate the ability to regulate the reorientation dynamics of active particles through mechanical changes in microstructured backgrounds.
@@
The manuscript is structured as follows. In Section 2, we first discuss the preparation of colloidal monolayers where the inter-particle interactions can be tuned using external AC electric fields. We then describe the two-dimensional motion of active particles throughout colloidal monolayers in term of their swimming velocity and persistence time. Here, the electric field determines not only the interactions between the passive particles of the monolayer, but also the propulsion speed of the active beads. In Section 3, we present a final discussion on our results. Sections 45 and 6 contain the experimental, numerical and theoretical protocols.
@@ 2.1 Tunable colloidal environments
[…]
@@
In simulations, we model the background matrix as a two-dimensional system of interacting Brownian particles, where hydrodynamics interactions are negligible and the solvent simply induces a thermal noise [47]. The particles interact via volume exclusion, modeled as a WCA potential [48], and a long-range dipole-dipole repulsion mimicking the dipolar interactions between $\rm SiO_{2}$ microparticles at $\rm E>0$. The strength of these potentials, $\epsilon$ and $\rm K$ respectively, is determined by matching the peaks of the pair correlation functions $\rm g(r)$ with experiments for every value of $\rm E$ and $\rm\phi$ (see Fig. 2(a), Fig. S1 and Table S1). Further experimental and numerical details are reported in Sec. 4 and Sec. 5, respectively.
Both in simulations and experiments, the long-range repulsive interactions induced by the external electric field increase the effective particle size and promote a shift of the freezing line of the two-dimensional phase diagram towards lower packing fractions. Features that are characteristic of solid structures (e.g. hexagonal crystalline domains) appear at smaller $\phi$, when the original microstructure (at $\rm E=0$) was in the fluid phase. The red curves in Fig. 2(a) show for example the pair correlation functions, $\rm g(r)$, for a monolayer at $\phi=0.48$ subjected to three different electric fields. As $\rm E$ increases, the peaks of $\rm g(r)$ become more pronounced until, at $\rm E=108$ $\rm V/mm$, the second peak between $\rm 4R_{p}$ and $\rm 6R_{p}$ splits, as reported for hexagonal structures. To quantify the appearance of solid domains, we consider the absolute value of the orientational hexagonal order parameter per particle:
@@
The structural changes induced by the electric field are observed for a broad range of packing fractions, suggesting the generality of the observed quenching mechanism. These results are shown in Fig. 2(f), where $\rm\left|\Psi_{6}\right|$ is reported as a function of the electric field $\rm E$ for four different values of $\phi$. The corresponding pair correlation functions $\rm g(r)$ and phases $\varphi$ of $\rm\Psi_{6,n}$ are shown in Fig. S1 and Fig. S2, respectively. In all instances, the global bond order parameter $\rm\left|\Psi_{6}\right|$ increases with $\rm E$, although we observe only a weak increase for large values of $\rm E$ at $\phi=0.15$ (green data) and a plateau in the densest environment ($\phi=0.64$, orange data). Indeed, in the small packing fraction regime, the smaller is the density, the larger is the increase of the dipole interaction strength needed to induce significant structural changes. By contrast, at high packing fractions, the observed plateau implies that the system has already reached the almost-close packing regime characterized by a defects-poor configuration. As a further confirmation, we illustrate in Fig. 2(g) the dynamical orientational correlation $\rm g_{6}(\Delta t)=\langle\Psi^{*}_{6,n}(\Delta t)\Psi_{6,n}(0)\rangle$ as a function of the delay time $\rm\Delta t/\tau_{B}$ normalized by Brownian diffusion time $\rm\tau_{B}=(6\pi\eta_{0}R_{p}^{3})/(k_{B}T)$, where $\eta_{0}$ is the water viscosity and $\rm k_{B}T$ is the thermal energy (see also Fig. S3 for a comparison with the numerical results). At $\phi=0.64$ (orange data), $\rm g_{6}(\Delta t)$ shows no significant decay within our experimental timescale, suggesting the presence of defect-free structures. At smaller packing fractions, $\rm g_{6}(\Delta t)$ decays at a rate that depends on both $\rm E$ and $\phi$. Interestingly, similar decays are recovered with different packing fractions and electric fields, in agreement with the measurements of $\rm\left|\Psi_{6}\right|$ (compare for instance the red and blue data in Fig. 2(f)). This demonstrates that the electric field can be used as a tool to create optimally ordered configurations without changing the packing fraction.
@@ 2.2 Active particles in tunable colloidal environments
Figure 3: Swimming in crowded environments. (a) Mean swimming velocities (black) $v_{0}$ and (red) $v$ plotted as a function of the electric-field strength squared $\rm E^{2}$ for active particles swimming “freely” ($\phi=0$) or cruising in colloidal environments of packing fraction $\phi=0.48$. The solid lines are linear fits crossing the origin. (b) Ratio between the effective viscosity $\eta_{\mathrm{eff}}$ of the colloidal environment and the water viscosity $\eta_{0}$, calculated as described in the text. (c–d) Normalized histograms of the instantaneous speed $|\dot{\mathbf{x}}|$ under an applied electric field: (c) $\rm E=42\,\mathrm{V/mm}$ and (d) $\rm E=108\,\mathrm{V/mm}$. (e–g) Trajectories of active particles cruising in colloidal environments of packing fraction $\phi=0.48$ under applied electric fields: (e) $\rm E=42\,\mathrm{V/mm}$, (f) $\rm E=75\,\mathrm{V/mm}$, and (g) $\rm E=108\,\mathrm{V/mm}$. The scale bar is dynamic and corresponds to $(v\tau_{\mathrm{R}})$ for all trajectories, where $v$ is the average swimming velocity and $\tau_{\mathrm{R}}$ is the rotational Brownian time (defined in the text). The color indicates the time: from $t=0$ (white) to $t=500\,\mathrm{s}$ (red). (h) Time autocorrelation functions of $\dot{\mathbf{x}}$ for the trajectories shown in (e–g). The solid lines are linear fits. (i) Mean persistence times $\tau$ (red symbols, $\phi=0.48$) and $\tau_{0}$ (black symbols, $\phi=0$) plotted against $v$. The filled symbols are obtained from experiments, while the empty symbols (linked by the dotted line) are from numerical simulations matching the experimental values of $v$. In all panels, the error bars represent the standard deviations.
@@
We then prepare the same colloidal environments as in Fig. 2 and add a very small number of self-propelling particles such that the interactions between them can be neglected. The active particles behave as ‘snowploughs’ making their way through the two-dimensional crowded environment (see Supplementary Videos S1 and S2, note that the SiO2 spheres do not get into physical contact). Even though they do not sense local defects and grain boundaries due to the large size (compared to the one of the bath passive particles), their active dynamics is strongly coupled to the monolayer’s properties (see Fig. 1a and Fig. 1b). In simulations, this behavior is reproduced through a 2D active Brownian particle dynamics where the active particle moves in the direction identified by the outward vector normal to the $\rm Au$ hemisphere. As for passive colloids, hydrodynamics interactions are neglected and the solvent induces translational and rotational noise on the particle position and orientation [52]. The active particle interacts with passive colloids via volume exclusion and repulsive dipolar forces. Volume exclusion is obtained from a WCA potential accounting for the larger size of the active colloid, whereas dipolar forces stem from the $\rm SiO_{2}$ back hemisphere of the active particle. As such, although active-passive dipolar interactions follow the $\rm~r^{-3}$ scaling (as in the case of uncoated passive particles), the force is not applied on the particle center. Instead, it acts the center of the SiO2 hemisphere, due to the fact that the gold cap and the passive particles do not exhibit dipolar mutual interactions. Hence, the orientational dynamics is not purely governed by rotational noise but it is additionally susceptible to local torques. In the following, we analyze the active trajectories by extracting the mean swimming velocity $v$ and persistence time $\tau$. The same quantities measured at $\phi=0$ (without the colloidal environment) are denoted as $v_{0}$ and $\tau_{0}$. As in the previous Section, we first describe the main results at $\phi=0.48$ and then show the generality of our findings at other packing fractions. Further information about the particle synthesis and self-propulsion method are in Sec. 4. The numerical simulations are described in Sec. 5.
@@
Nonetheless, a description based on an average swimming velocity of the self-propelling particles and an effective viscosity of the colloidal environment does not take into account the microstructure of the monolayer. A closer look at the instantaneous velocities sometimes reveals large fluctuations around their mean value, where $\left|\dot{\mathbf{x}}\right|$ drops to nearly zero before increasing again. This local and transient caging is due to the microstructure of the surrounding monolayer as well as the strength of interaction between its building blocks [29, 53]; it is qualitatively illustrated in the Supplementary Video S6 and pinpointed in Fig. 3(c-d) by plotting the probability distribution ${\rm p}(\left|\dot{\mathbf{x}}\right|)$. For small applied electric fields ($\rm E=42$ $\rm V/mm$, Fig. 3(c)), the dipolar interactions are weak and the active particle can navigate through the environment owing to the high mobility of the uncoated $\rm SiO_{2}$ microspheres. The corresponding distribution $\rm p(\left|\dot{\mathbf{x}}\right|)$ is therefore similar to those of free active particles (see also Figure S5, gray histograms). At higher electric fields ($\rm E=108$ $\rm V/mm$, Fig. 3(d)), $\rm p(\left|\dot{\mathbf{x}}\right|)$ shows a peak at $\left|\dot{\mathbf{x}}\right|\sim 0$ followed by a distribution at higher speeds; the active motion becomes intermittent and particles are occasionally caged by the surrounding microstructure. Velocity distributions for intermediate values of $\rm E$ are reported in Fig. S5 and show a consistent transition towards a bimodal distribution, as $\rm E$ increases. We report no caging at smaller packing fractions ($\phi=0.35$ and $\phi=0.15$). The same qualitative behaviour is also observed in numerical simulations (Fig. S5), although caging occurs therein more often. The discrepancy is likely due to the fact that the numerical system is perfectly two dimensional, whereas in experiments the active particles protrude above the colloidal monolayer because of their larger size (see Fig. 5 in Sec. 4). To support this hypothesis, Supplementary Video S7 shows an active particle being frequently caged in a monolayer of passive spheres of similar size.
@@
An increase of effective viscosity due to larger $\phi$ or $\rm E$ should also imply a slowdown of rotational Brownian motion of the active particles in the $xy$-plane and, consequently, a higher persistence time of their directed motion (i.e. a larger $\tau$). The data in Figure 3 suggest the opposite. Fig. 3(e-g) shows three typical active trajectories at fixed $\phi=0.48$ and different applied electric fields. As $\rm E$ increases from $42$ $\rm V/mm$ (Fig. 3(e)) to $108$ $\rm V/mm$ (Fig. 3(g)), the particles not only become faster, but also change their swimming direction more frequently, as shown by comparing the typical straight path of each trajectories to the persistence length $(v\tau_{\rm R})$ expected as a result of free rotational diffusion with timescale $\rm\tau_{R}$. We extract the characteristic persistence times $\tau$ and $\tau_{0}$ from the linear fit (Fig. 3(h), black lines) of the time autocorrelation function of the instantaneous velocity vector $\rm\left\langle\dot{\mathbf{x}}(\Delta t)\cdot\dot{\mathbf{x}}(0)\right\rangle$ (Figure 3(h), red circles), as described in Sec. 4. Faster decays indicate smaller persistence times. At $\phi=0.48$, the velocity autocorrelation decays faster for larger applied electric fields. This corresponds to mean persistence times that become more than one order of magnitude smaller as $\rm E$ goes from $\rm E=42$ $\rm V/mm$ to $\rm E=108$ $\rm V/mm$, i.e. as the mean swimming velocity increases (Fig. 3(i), red symbols). This is in stark contrast to the motion of ‘free’ active particles ($\phi=0$) where $\tau_{0}$ remains approximately constant at all applied electric fields and swimming velocities (see black symbols in Fig. 3(i) as well as Fig. S4). The measured persistence times are also in agreement with numerical simulations (Fig. 3(i), empty red symbols) that are set to match the mean swimming velocity $v$ of the active particles and the pair correlation function $\rm g(r)$ of the colloidal bath.